# Verina: Benchmarking Verifiable Code Generation

### v1

@@ 4 True
Specification generation failures: unsound post-conditions. Figure 17 illustrates how LLMs can generate post-conditions that miss critical constraints, leading to unsound specifications. The task involves adding two numbers represented as digit lists in reverse order. The ground truth post-condition correctly enforces three essential properties: arithmetic correctness, digit validity (each digit < 10), and prohibition of leading zeros except for the special case where the result is zero. However, the LLM-generated post-condition omits the leading zero constraint entirely, only checking that the result is non-empty and digits are valid. This unsound specification would accept invalid outputs like [2, 1, 0] representing 012 (which is one of the negative test case), which should be rejected in favor of [2, 1]. The missing constraint demonstrates how LLMs may capture the primary functional requirements while overlooking more subtle constraints that are crucial for correctness.
@@
Figure 17: Example (verina_advanced_5): o4-mini generates unsound postcondition that fails to rule out leading zeros in the result.
@@
Figure 18: Example (verina_advanced_72): o4-mini generates unsound and incomplete post-condition that fails to handle the edge case n = 0.
@@
Figure 19: Example (verina_advanced_25): o4-mini generates post-condition with quantifiers over lists that cannot be evaluated by plausible testing.
Proof generation success with iterative refinement. Figure 20, 21, 22, and 23 demonstrate o4-mini’s iterative proof refinement process over 24 attempts. The task involves proving that a list element removal function satisfies its specification. Initially, o4-mini generates an overly simple proof using direct induction and simp tactics, which fails to handle the complex logical structure. After 23 failed refinement attempts, the model makes useful improvements based on Lean compiler error messages, significantly restructuring its approach. The error message from the 23rd attempt reveals that the proof got stuck on conditional expressions involving decide predicates. In the successful 24th attempt, o4-mini improves its strategy by introducing an auxiliary lemma helper_eq_filter and using explicit case analysis with by_cases. This demonstrates how iterative refinement with compiler feedback can guide LLMs toward substantially different and correct proof strategies, though it may require many attempts to achieve success.
@@
Figure 20: Example (verina_advanced_64): o4-mini generates incorrect proof on the first attempt.
@@
Figure 21: Example (verina_advanced_64): o4-mini generates incorrect proof after 23 refinement attempts.
@@
Figure 22: Example (verina_advanced_64): Lean 4 compiler error message from the 23rd refinement attempt.
@@
Figure 23: Example (verina_advanced_64): o4-mini generates correct proof on the 24th attempt.

### v2

@@ 4 True
Specification generation failures: unsound post-conditions. Figure 19 illustrates how LLMs can generate post-conditions that miss critical constraints, leading to unsound specifications. The task involves adding two numbers represented as digit lists in reverse order. The ground truth post-condition correctly enforces three essential properties: arithmetic correctness, digit validity (all digits should be less than 10), and prohibition of leading zeros except for the special case where the result is zero. However, the LLM-generated post-condition omits the leading zero constraint entirely, only checking that the result is non-empty and digits are valid. This unsound specification would accept invalid outputs like [2, 1, 0] representing 012 (which is one of the negative test case), which should be rejected in favor of [2, 1]. The missing constraint demonstrates how LLMs may capture the primary functional requirements while overlooking more subtle constraints that are crucial for correctness.
@@
Figure 19: Example (verina_b_5): o4-mini generates unsound postcondition that fails to rule out leading zeros in the result.
@@
Figure 20: Example (verina_b_72): o4-mini generates unsound and incomplete post-condition that fails to handle the edge case n = 0.
@@
Figure 21: Example (verina_b_25): o4-mini generates post-condition with quantifiers over lists that cannot be evaluated by plausible testing.
Proof generation success with iterative refinement. Figure 22, 23, 24, and 25 demonstrate o4-mini’s iterative proof refinement process over 24 attempts. The task involves proving that a list element removal function satisfies its specification. Initially, o4-mini generates an overly simple proof using direct induction and simp tactics, which fails to handle the complex logical structure. After 23 failed refinement attempts, the model makes useful improvements based on Lean compiler error messages, significantly restructuring its approach. The error message from the 23rd attempt reveals that the proof got stuck on conditional expressions involving decide predicates. In the successful 24th attempt, o4-mini improves its strategy by introducing an auxiliary lemma helper_eq_filter and using explicit case analysis with by_cases. This demonstrates how iterative refinement with compiler feedback can guide LLMs toward substantially different and correct proof strategies, though it may require many attempts to achieve success.
@@
Figure 22: Example (verina_b_64): o4-mini generates incorrect proof on the first attempt.
@@
Figure 23: Example (verina_b_64): o4-mini generates incorrect proof after 23 refinement attempts.
@@
Figure 24: Example (verina_b_64): Lean 4 compiler error message from the 23rd refinement attempt.
@@
Figure 25: Example (verina_b_64): o4-mini generates correct proof on the 24th attempt.
Appendix D Use of LLM
LLMs were used in a limited and supervised manner during the construction of Verina. Specifically, we employed OpenAI o3-mini with few-shot prompting to assist in translating 59 Dafny instances from CloverBench (Sun et al., 2024) into Lean, as discussed in Section 3.2. All such translations were subsequently inspected, corrected, and verified by the authors to ensure accuracy. In addition, LLMs were used as assistive tools for editing and polishing the presentation of the paper. LLMs were not involved in research ideation, discovery of related work, experimental design, dataset selection, or analysis.